1. An executive nearing retirement made two investments totaling $15,000. In one year, these investments yielded $1432 in simple interest. Part of the money was invested at 9% and the rest at 10%. How much was invested at each rate?
2. Two cars leave Salt Lake City, travelng in opoosite directions. One car travels at a pseed of 80 km/h and the other at 96 hm/h. In how many hours will they 528 km apart?
3. McDonald's recently sold small soft drinks for $.87 meidum soft drinks for $1.08, and large soft drinks for $1.54. During a lunchtime rush, Chris sold 40 soft dirnks for a total of $43.40. The number of small and large drinks, combined, was 10 fewer than the number of medium drinks. How many drinks of each size were sold?

Answer by stanbon(48554)

1. An executive nearing retirement made two investments totaling $15,000. In one year, these investments yielded $1432 in simple interest. Part of the money was invested at 9% and the rest at 10%. How much was invested at each rate?
Let amt. invested at 9% be "x"; Interest on this is 0.09x dollars
Amt. invested at 10% is "15000-x" ; Interest on this is 0.1(15000-x)=1500-0.1x
dollars
EQUATION:
interext + interest=1432 dollars
0.09x + 1500-0.10x = 1432
-0.01x = -68
x=$6800 (amt invested at 9%)
15000-6800=$8200 (amt invested at 10%
--------------------------------
2. Two cars leave Salt Lake City, travelng in opoosite directions. One car travels at a pseed of 80 km/h and the other at 96 hm/h. In how many hours will they 528 km apart?
let "x" be the number of hours before they meet.
EQUATION:
80x + 96x = 528
178x=528
x=3 hours
---------------------------
3. McDonald's recently sold small soft drinks for $.87 meidum soft drinks for $1.08, and large soft drinks for $1.54. During a lunchtime rush, Chris sold 40 soft dirnks for a total of $43.40. The number of small and large drinks, combined, was 10 fewer than the number of medium drinks. How many drinks of each size were sold?
Let S,M,L be the number of small, medium, and large drinks.
S+M+L=40
87S+108M+154L=4340
S-M+L=-10
==============
Solve this system of three equations.
I used a matrix method and got:
S=10
M=25
L=5
-----------
Cheers,
Stan H.